OpenStudy (anonymous):

appreciate the help! Solve using substitution 8x-7y=-5 x-8y=-2

OpenStudy (anonymous):

second equation solve for x via \[x=8y-2\]

OpenStudy (anonymous):

8(8y-2)-7y ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

don't forget the =-5, gets important

OpenStudy (anonymous):

then replace in first equation being careful with parentheses \[8(8y-2)-7y=-5\] \[64y-16-7y=-5\] \[48y=-5+16\] \[48y=11\]

OpenStudy (anonymous):

oh mistake sorry

OpenStudy (anonymous):

64-7=57

OpenStudy (anonymous):

y=11/57 and subsitute in

OpenStudy (anonymous):

\[8(8y-2)-7y=-5\] \[64y-16-7y=-5\] \[57y=-5+16\] \[57y=11\] \[y=\frac{11}{57}\]

OpenStudy (anonymous):

looks like you got it!

OpenStudy (anonymous):

^what he said

OpenStudy (anonymous):

x=-26?

OpenStudy (anonymous):

x=(88/57)-(114/57) x=-26/57